The Hypotenuse of the right triangle is 15 centimeters long. One leg is 9 centimeters long. How long is the other leg?
Answer by waynest(281)

hypotenuse^2 = a^2 + b^2
hypotenuse = 15
a = 9
b = ?
15^2 = 9^2 + b^2
15^2 - 9^2 = b^2
225 - 81 = b^2
144 = b^2
sqrt144 = sqrtb^2
12 = b